OpenStudy (anonymous):

For the equation x2 + 3x + j = 0, find all the values of j such that the equation has two real number solutions.

OpenStudy (phi):

Do you know what the discriminant is? It must be 0 or positive for pure real roots.

OpenStudy (anonymous):

thats all they gave me for the problem

OpenStudy (phi):

Here are the gory details http://en.wikipedia.org/wiki/Quadratic_equation#Quadratic_formula

OpenStudy (anonymous):

I don't get how to work this out

OpenStudy (anonymous):

do you by chance have any advice on how to work it out?

OpenStudy (phi):

You match your equation to the form \[a x^2 +bx+c =0\] \[ x^2 + 3x + j = 0 \] so what is a? b? c? once you know a,b,c, you use them in \[ b^2 -4ac ≥ 0\]

OpenStudy (phi):

For example, in \[ x^2 +3x +j=0 \] a must be 1